OpenStudy (anonymous):

-2/3 sqrt63x^3 Simplify :

OpenStudy (anonymous):

\[-\frac{ 2 }{ 3 }\sqrt{63x^3}\] does this look right?

OpenStudy (anonymous):

\[-\frac{ 2 }{ 3}\sqrt{3^2*7*x^2*x}\] realize what can be factored out as squares, then pull them out of the square root

OpenStudy (anonymous):

\[-\frac{ 2 }{ 3 }*3*x \sqrt{7x}\]